# M4-SAR: A Multi-Resolution, Multi-Polarization, Multi-Scene, Multi-Source Dataset and Benchmark for optical-SAR Fusion Object Detection

### v2

@@ Abstract
Single-source remote sensing object detection using optical or SAR images struggles in complex environments. Optical images offer rich textural details but are often affected by low-light, cloud-obscured, or low-resolution conditions, reducing the detection performance. SAR images are robust to weather, but suffer from speckle noise and limited semantic expressiveness. Optical and SAR images provide complementary advantages, and fusing them can significantly improve the detection accuracy. However, progress in this field is hindered by the lack of large-scale, standardized datasets. To address these challenges, we propose a new comprehensive dataset for optical-SAR fusion object detection, named Multi-resolution, Multi-polarization, Multi-scene, Multi-source SAR dataset (M4-SAR). It contains 112,174 precisely aligned image pairs and nearly one million labeled instances with arbitrary orientations, spanning six key categories. To enable standardized evaluation, we develop a unified benchmarking toolkit that integrates six state-of-the-art multi-source fusion methods. Additionally, we propose E2E-OSDet, a novel end-to-end multi-source fusion detection framework that mitigates cross-domain discrepancies and establishes a robust baseline for future studies. Extensive experiments on M4-SAR demonstrate that fusing optical and SAR data can improve mAP by 5.7% over single-source inputs, with particularly significant gains in complex environments. The dataset and code are publicly available at https://github.com/wchao0601/M4-SAR.
I Introduction
@@
With the rapid development of deep learning techniques, single-source detection (e.g., optical [24, 31] or synthetic aperture radar (SAR) [47]), and multi-source detection (optical-SAR [23, 62]) have garnered significant attention in the remote sensing field. These methods have demonstrated remarkable effectiveness in applications, including disaster monitoring and urban planning. As shown in Fig. 1 (c), optical images provide high-resolution visual cues with rich texture, color, and spectral information, facilitating fine-grained feature extraction and semantic interpretation under clear atmospheric conditions. However, their detection performance degrades markedly under low-light, cloud-obscured, or low-resolution conditions (see Fig. 1 (a)). Despite these drawbacks, the intuitive appearance and broad availability of optical data have fueled significant advancements in optical image object detection [28, 68, 3, 5]. In contrast, SAR sensors enable all-weather, all-day imaging, resilient to cloud cover and illumination changes. However, the inherent speckle noise and low contrast arising from microwave backscatter present significant obstacles to precise localization and classification. Moreover, constructing large-scale SAR datasets [25, 67, 54] is hindered by the high cost of data acquisition and the laborious fine-grained annotation. Although optical and SAR images exhibit strong complementarity in spatial resolution and environmental robustness, existing single-source [43, 65, 41, 40] and multi-source approaches [14, 2] face challenges in addressing cross-domain distribution shifts and suffer from the absence of a standardized evaluation benchmark. To enable robust detection in complex scenarios, there is an urgent need for a large-scale, precisely aligned optical-SAR fusion detection dataset and a standardized benchmarking toolkit.
Currently, a major bottleneck in optical-SAR fusion object detection lies in the lack of large-scale, high-quality, and standardized datasets. As shown in Table I, most existing SAR datasets are limited to single-source data, suffering from limited scale, narrow category coverage, or inconsistent annotations. For example, HRSID [53], SADD [64], SSDD [66], SIVED [26], and ShipDataset [52] exhibit notable limitations in terms of category diversity and dataset scale. In recent years, the emergence of larger-scale datasets such as SARDet-100K [25], FAIR-CSAR [54], and RSAR [67] has partially addressed these limitations. However, these datasets are often synthesized by standardization of data from multiple sources, potentially introducing bias during data preprocessing and compromising the fairness of model evaluation. Moreover, constructing paired optical-SAR datasets is prohibitively expensive: acquiring SAR data is already costly, and the acquisition of spatio-temporally aligned optical images is even more difficult, further limiting the construction of multi-source datasets.
Although OGSOD-1.0 [49] and OGSOD-2.0 [48] represent pioneering efforts in optical-SAR multi-source data, they are primarily designed for cross-modal knowledge distillation. Moreover, their test sets consist solely of SAR images, which limits their applicability for end-to-end multi-source fusion object detection. To address this issue, we constructed the novel optical-SAR fusion object detection dataset (M4-SAR) covering multi-resolution, multi-polarization, multi-scene, and multi-source using publicly available optical and SAR data provided by Sentinel-1 [44], and Sentinel-2 [7] satellites. To overcome SAR annotation challenges, we propose a semi-supervised optical-assisted labeling strategy that utilizes optical images’ semantic richness to substantially improve annotation quality. As illustrated in Fig. 1 and Table I, M4-SAR includes six key target categories with 112,174 image pairs and 981,862 instances. M4-SAR surpasses existing datasets in instance volume and alignment precision, establishing a robust benchmark for multi-source fusion in complex scenarios.
@@
Another important challenge is the lack of publicly available algorithms tailored for optical-SAR fusion object detection. Most existing studies are based on single-source datasets or evaluated on limited scenarios, making it difficult to conduct fair performance comparisons. To bridge this gap, we introduce Multi-Source Rotated Object Detector (MSRODet), an open-source toolkit that integrates state-of-the-art fusion detectors (e.g., CFT [36], CLANet [14], CSSA [2], CMADet [39], ICAFusion [38], and MMIDet [60]), providing a standardized evaluation framework for researchers utilizing the M4-SAR dataset. MSRODet not only enables fair and reproducible benchmarking across different fusion methods but also lays a solid foundation for the development of future multi-source fusion object detection algorithms.
Moreover, extensive experiments demonstrate that the intrinsic differences in the imaging mechanisms of optical and SAR sensors lead to significant domain gaps (Fig. 4 (a)), thereby constraining the performance of optical-SAR fusion object detection. To address this issue, we propose the first end-to-end framework for optical-SAR fusion object detection, termed E2E-OSDet. This framework effectively fuses features from both modalities, alleviates cross-modal domain gaps, and results in significant improvements in both inference speed and detection performance.
This work offers substantial contributions in terms of dataset construction, toolkit development, and algorithm design. The main contributions of this work are summarized as follows:
  • M4-SAR Dataset: M4-SAR is a novel optical-SAR fusion object detection dataset with multi-resolution, multi-polarization, and multi-scene characteristics, containing 112,174 image pairs and 981,862 instances. This dataset comprehensively covers complex scenes and diverse target categories frequently encountered in multi-source fusion detection tasks, thereby providing a robust foundation for advancing research in this field.
  • Multi-source Evaluation Toolkit (MSRODet) and E2E-OSDet: We introduce MSRODet, a comprehensive benchmarking toolkit incorporating state-of-the-art fusion methods tailored for optical-SAR object detection, which provides a standardized and reproducible evaluation platform based on the M4-SAR dataset. In addition, we propose E2E-OSDet, a novel end-to-end optical-SAR fusion object detection framework, which effectively mitigates significant cross-domain discrepancies between optical and SAR modalities at the feature level.
@@
TABLE I: Comparison between existing optical and SAR object detection datasets and the proposed M4-SAR dataset.
  Dataset OPT SAR Spatial Resolution Polarization Image Size Category Instances Images Ins/Img
HRSC2016 [28] $\checkmark$ 0.4M $\sim$ 2M - 300 $\sim$ 1,500 1 2,976 1,061 2.80
LEVIR [68] $\checkmark$ 0.2M $\sim$ 1M - 600 $\times$ 800 3 11,028 21,952 0.50
DIOR-R [3] $\checkmark$ 0.5M $\sim$ 30M - 800 $\times$ 800 20 190,288 23,463 8.11
DOTA-v1.0 [5] $\checkmark$ - - 800 $\sim$ 4,000 15 188,282 2,806 67.10
HRSID [53] $\checkmark$ - HH, HV, VH, VV 800 $\times$ 800 1 16,969 5,604 3.03
SADD [64] $\checkmark$ - $H$ 224 $\times$ 224 1 7,835 883 8.87
SSDD [66] $\checkmark$ - HH, HV, VH, VV 512 $\times$ 512 1 2,587 1,160 2.23
SIVED [26] $\checkmark$ - VV, VH 512 $\times$ 512 1 12,013 1,044 11.51
ShipDataset [52] $\checkmark$ - HH, HV, VH, VV 256 $\times$ 256 1 50,885 39,729 1.28
MSAR [55] $\checkmark$ - HH, HV, VH, VV 256 $\times$ 256 4 65,202 30,158 2.16
SARDet-100K [25] $\checkmark$ - HH, HV, VH, VV 512 $\times$ 512 6 245,653 116,598 2.11
FAIR-CSAR [54] $\checkmark$ - HH, HV, VH, VV 1,024 $\times$ 1,024 5 349,002 106,672 2.11
RSAR [67] $\checkmark$ - HH, HV, VH, VV 512 $\times$ 512 6 183,534 95,842 1.91
OGSOD-1.0 [49] $\checkmark$ $\checkmark$ 20M VV, VH 256 $\times$ 256 3 48,589 18,331 2.65
OGSOD-2.0 [48] $\checkmark$ $\checkmark$ 20M VV, VH 256 $\times$ 256 3 54,000 20,359 2.65
M4-SAR (Ours) $\checkmark$ $\checkmark$ 10M, 60M VV, VH 512 $\times$ 512 6 981,862 112,174 8.75
II Related Work
II-A Related Datasets
Progress in SAR object detection is impeded by the high costs of data acquisition and the intricacy of fine-grained annotation. Consequently, existing datasets, including HRSID [53], SADD [64], SSDD [66], and SIVED [26], are frequently restricted to a limited number of common categories (e.g., ships), generally comprising fewer than 10,000 images with constrained class diversity (Table I). These constraints impede the development of robust and generalizable models for remote sensing applications. Recent efforts have sought to tackle these challenges. Wang et al. [52] introduced the Ship-Dataset, which comprises 39,729 ship patches extracted from Gaofen-3 and Sentinel-1 imagery, extensive in scale but constrained in semantic diversity. Xia et al. [55] introduced MSAR, broadening the categories to include airplanes, tanks, bridges, and ships. To address data scarcity, Li et al. [25] released SARDet-100K, comprising 116,598 images and 245,653 instances across six categories. Wu et al. [54] contributed FAIR-CSAR, providing fine-grained multi-domain annotations, while Zhang et al. [67] enhanced annotation efficiency in RSAR through weakly supervised learning and rotated bounding boxes.
Despite these advancements, which signify a shift from small-scale, single-category collections to large-scale, multi-class benchmarks, most existing datasets remain restricted to single-source data. Consequently, they cannot exploit complementary multimodal information critical for robust object detection in real-world remote sensing scenarios. Although OGSOD-1.0 [49] and OGSOD-2.0 [48] offer aligned optical-SAR image pairs, their evaluation protocol depends solely on SAR data, limiting their applicability for end-to-end fusion-based detection. Thus, there is an urgent need for a high-quality multi-source dataset that enables comprehensive optical-SAR fusion object detection.
II-B Remote Sensing Object Detection
Remote sensing images, encompassing both optical and synthetic aperture radar (SAR) modalities, pose distinct challenges for object detection. SAR imagery is particularly susceptible to multiplicative speckle noise and structural distortions, which hinder reliable feature extraction. Early approaches sought to address these challenges using handcrafted descriptors, including the Canny edge detector [1], Wavelet Scattering Transform (WST) [34], Haar-like features [45], Histogram of Oriented Gradients (HOG) [4], and Gradient-by-Ratio edge detection (Grad) [21]. These methods capture structural information through edge detection, frequency decomposition, or gradient statistics, and have been effectively adapted for both optical and SAR targets. However, their constrained representation capacity and susceptibility to imaging conditions limit their performance in complex remote sensing environments.
With the advent of deep learning, object detection has attained remarkable advancements. Architectures such as R-CNN [9, 12], YOLO [16, 17, 50, 46], and DETR have exhibited robust generalization across natural images, with numerous variants adapted for remote sensing data. Nevertheless, SAR object detection remains formidable due to speckle noise and the absence of rich semantic cues. To tackle these challenges, recent studies have developed SAR-specific designs. For instance, the SFS-Conv [19] employs fractional Gabor transforms to enhance frequency-domain representations, while SAR-JEPA [21] introduces a joint embedding framework that predicts multi-scale gradient features through masked local reconstruction. The large-margin anomaly detection model [18] facilitates cross-modal generalization between optical and SAR data by learning a deviation metric independent of the background distribution. Wang et al. developed a lightweight SAR ship detection network through evolutionary strategies [51], demonstrating the potential of resource-efficient architectures. Beyond these task-specific advancements, foundational SAR understanding models such as SARATR-X [20] reveal that self-supervised pretraining on large-scale SAR datasets produces highly transferable feature representations.
II-C Multi-source Fusion Object Detection
Deep neural networks, due to their strong nonlinear modeling capacity, have significantly advanced multi–source fusion object detection [39, 2, 38]. Recent studies have tackled key challenges in cross-modal integration. For example, Yuan et al. [59] and Fang et al. [36] proposed alignment strategies to mitigate spatial misalignments between heterogeneous modalities. He et al. [14] investigated semantic conflicts and complementarity across sources to enhance fusion effectiveness. Zeng et al. [60] introduced a contour enhancement module to extract modality-specific spectral features from visible and infrared images, emphasizing spatial diversity and fine-grained structures. Despite these advances, many existing frameworks suffer from complex training procedures, limiting scalability and real-world deployment. To address this, Zhang et al. [63] proposed E2E-MFD, an end-to-end architecture that streamlines the fusion pipeline through a single-stage training process while achieving strong performance.
@@ II-C Multi-source Fusion Object Detection
III M4-SAR Dataset
III-A Overview of the Proposed Dataset Construction Pipeline
@@
Fig. 2 illustrates the overall pipeline of the proposed dataset construction process. In the first stage, we acquire optical and SAR images from Sentinel-1 and Sentinel-2 satellites provided by the European Space Agency (ESA). In the second stage, the acquired images are pre-processed and aligned. After completing the alignment, we crop aligned image pairs into fixed-size patches to generate initial unlabeled samples. In the third stage, we introduce a semi-supervised annotation strategy guided by optical images to ensure high-quality labeling of the image pairs.
III-B Data Collection Procedure
@@
First, we collected the required data via the open-access platform provided by the ESA. A map-based selection tool was used to specify areas of interest, covering representative coastal cities, including Tianjin, Los Angeles, Hong Kong, and London. We then retrieved SAR images along with their corresponding digital elevation model (DEM) data according to specific criteria. For optical imagery, Sentinel-2 images covering the same regions as the Sentinel-1 data were acquired, specifically including the red (R), green (G), and blue (B) spectral bands. To minimize variations caused by temporal discrepancies between the two modalities, we selected image pairs with a time interval of less than 10 days. This setting ensures that targets of interest (e.g., bridges, oil tanks) maintain structural consistency within the short interval, thereby enhancing the accuracy of subsequent instance-level alignment. The product type for optical images is L2A, with cloud coverage limited to 0–10% (the corresponding cloud-free (0%) images are used for high-quality annotation) to guarantee high visual fidelity. The SAR images are provided in the Single Look Complex (SLC) format to preserve full coherence information. The collected data span the period from 2020 to 2022 and encompass multiple spatial resolutions and polarization modes. The optical images have 10m, 20m, and 60m resolutions, and the SAR data cover both vertical-horizontal (VH) and vertical-vertical (VV) polarizations.
III-C Data Processing

### v3

# https://github.com/wchao0601/M4-SAR

@@ Abstract
Single-source remote sensing object detection using optical or SAR images struggles in complex environments. Optical images offer rich textural details but are often affected by low-light, cloud-obscured, or low-resolution conditions, reducing the detection performance. SAR images are robust to weather, but suffer from speckle noise and limited semantic expressiveness. Optical and SAR images provide complementary advantages, and fusing them can significantly improve the detection accuracy. However, progress in this field is hindered by the lack of large-scale, standardized datasets. To address these challenges, we propose a new comprehensive dataset for optical-SAR fusion object detection, named Multi-resolution, Multi-polarization, Multi-scene, Multi-source SAR dataset (M4-SAR). It contains 112,174 instance-level aligned image pairs and nearly one million labeled instances with arbitrary orientations, spanning six key categories. To enable standardized evaluation, we develop a unified benchmarking toolkit that integrates six state-of-the-art multi-source fusion methods. Additionally, we propose E2E-OSDet, a novel end-to-end multi-source fusion detection framework that mitigates cross-domain discrepancies and establishes a robust baseline for future studies. Extensive experiments on M4-SAR demonstrate that fusing optical and SAR data can improve mAP by 5.7% over single-source inputs, with particularly significant gains in complex environments.
footnotetext: Corresponding Author
1 Introduction
@@
With the rapid development of deep learning techniques, single-source object detection (e.g., optical [li2024lsknet, lu2025lwganet, lu2026unravelnet] or synthetic aperture radar (SAR) [wang2025msod, wang2026localized]), and multi-source object detection (optical-SAR [li2021deep, zhang2022domain]) have garnered significant attention in the remote sensing field. These methods have demonstrated remarkable effectiveness in applications, including disaster monitoring and urban planning. As shown in Fig. 1 (c), optical images provide high-resolution visual cues with rich texture, color, and spectral information, facilitating fine-grained feature extraction and semantic interpretation under clear atmospheric conditions. However, their detection performance degrades markedly under low-light, cloud-obscured, or low-resolution conditions (see Fig. 1 (a)). Despite these drawbacks, the intuitive appearance and broad availability of optical data have fueled significant advancements in optical image object detection [liu2017high, zou2017random, cheng2022anchor, ding2021object]. In contrast, SAR sensors enable all-weather, all-day imaging, resilient to cloud cover and illumination changes. However, the inherent speckle noise and low contrast arising from microwave backscatter present significant obstacles to precise localization and classification. Moreover, constructing large-scale SAR datasets [li2024sardet, zhang2025rsar, wu2024fair] is hindered by the high cost of data acquisition and the laborious fine-grained annotation. Although optical and SAR images exhibit strong complementarity in spatial resolution and environmental robustness, existing single-source [tian2019fcos, zhang2020bridging, sun2024united, WEFT] and multi-source object detection methods [he2023multispectral, cao2023multimodal] face challenges in addressing cross-domain distribution shifts and suffer from the absence of a standardized evaluation benchmark. To enable robust detection in complex scenarios, there is an urgent need for a large-scale optical-SAR fusion object detection dataset and a standardized benchmarking toolkit.
@@
Currently, a major bottleneck in optical-SAR fusion object detection lies in the lack of large-scale, high-quality, and standardized datasets. As shown in Table 1, most existing SAR datasets are limited to single-source data, suffering from limited scale, narrow category coverage, or inconsistent annotations. For example, HRSID [wei2020hrsid], SADD [zhang2022sefepnet], SSDD [zhang2021sar], SIVED [lin2023sived], and ShipDataset [wang2019sar] exhibit notable limitations in terms of category diversity and dataset scale. In recent years, the emergence of larger-scale datasets such as SARDet-100K [li2024sardet], FAIR-CSAR [wu2024fair], and RSAR [zhang2025rsar] has partially addressed these limitations. However, these datasets are often synthesized by standardization of data from multiple sources, potentially introducing bias during data preprocessing and compromising the fairness of model evaluation. Moreover, constructing paired optical-SAR datasets is prohibitively expensive: acquiring SAR data is already costly, and the acquisition of spatio-temporally aligned optical images is even more difficult, further limiting the construction of multi-source datasets.
Although OGSOD-1.0 [wang2023category] and OGSOD-2.0 [wang2025cross] represent pioneering efforts in optical-SAR multi-source data, they are primarily designed for cross-modal knowledge distillation. Moreover, their test sets consist solely of SAR images, which limits their applicability for end-to-end multi-source fusion object detection (Other related work can be found in Appendix 8.). To address this issue, we constructed the novel optical-SAR fusion object detection dataset (M4-SAR) covering multi-resolution, multi-polarization, multi-scene, and multi-source using publicly available optical and SAR data provided by Sentinel-1 [torres2012gmes], and Sentinel-2 [drusch2012sentinel] satellites. To overcome SAR annotation challenges, we propose a semi-supervised optical-assisted labeling strategy that utilizes optical images’ semantic richness to substantially improve annotation quality. As illustrated in Fig. 1 and Table 1, M4-SAR includes six key target categories with 112,174 image pairs and 981,862 instances. M4-SAR surpasses existing datasets in instance volume and image numbers, establishing a robust benchmark for multi-source fusion object detection in complex scenarios.
@@
Table 1: Comparison between existing optical and SAR object detection datasets and the proposed M4-SAR dataset. O and S denote optical and SAR data, respectively.
  Dataset O S Resolution Image Size Cat. Instances Images Ins/Img
HRSC2016 [liu2017high] $\checkmark$ 0.4M $\sim$ 2M 300 $\sim$ 1,500 1 2,976 1,061 2.80
[…]
DIOR-R [cheng2022anchor] $\checkmark$ 0.5M $\sim$ 30M $\times800$ 20 190,288 23,463 8.11
DOTA-v1.0 [ding2021object] $\checkmark$ - 800 $\sim$ 4,000 15 188,282 2,806 67.10
HRSID [wei2020hrsid] $\checkmark$ - $\times800$ 1 16,969 5,604 3.03
SADD [zhang2022sefepnet] $\checkmark$ - $\times224$ 1 7,835 883 8.87
SSDD [zhang2021sar] $\checkmark$ - $\times512$ 1 2,587 1,160 2.23
SIVED [lin2023sived] $\checkmark$ - $\times512$ 1 12,013 1,044 11.51
ShipDataset [wang2019sar] $\checkmark$ - $\times256$ 1 50,885 39,729 1.28
MSAR [xia2022crtranssar] $\checkmark$ - $\times256$ 4 65,202 30,158 2.16
SARDet-100K [li2024sardet] $\checkmark$ - $\times512$ 6 245,653 116,598 2.11
FAIR-CSAR [wu2024fair] $\checkmark$ - $\times1,024$ 5 349,002 106,672 2.11
RSAR [zhang2025rsar] $\checkmark$ - $\times512$ 6 183,534 95,842 1.91
OGSOD-1.0 [wang2023category] $\checkmark$ $\checkmark$ 20M $\times256$ 3 48,589 18,331 2.65
OGSOD-2.0 [wang2025cross] $\checkmark$ $\checkmark$ 20M $\times256$ 3 54,000 20,359 2.65
M4-SAR (Ours) $\checkmark$ $\checkmark$ 10M, 60M $\times512$ 6 981,862 112,174 8.75
Another important challenge is the lack of publicly available algorithms tailored for optical-SAR fusion object detection. Most existing studies are based on single-source datasets or evaluated on limited scenarios, making it difficult to conduct fair performance comparisons. To bridge this gap, we introduce Multi-Source Rotated Object Detector (MSRODet), an open-source toolkit that integrates state-of-the-art fusion detectors (e.g., MHFNet [zhang2026mhfnet], CFT [qingyun2022cross], CLANet [he2023multispectral], CSSA [cao2023multimodal], CMADet [song2024misaligned], ICAFusion [shen2024icafusion], and MMIDet [zeng2024mmi]), providing a standardized evaluation framework for researchers utilizing the M4-SAR dataset. MSRODet not only enables fair and reproducible benchmarking across different fusion methods but also lays a solid foundation for the development of future multi-source fusion object detection algorithms.
Moreover, extensive experiments demonstrate that the intrinsic differences in the imaging mechanisms of optical and SAR sensors lead to significant domain gaps (Fig. 4 (a)), thereby constraining the performance of optical-SAR fusion object detection. To address this issue, we propose the first end-to-end framework for optical-SAR fusion object detection, termed E2E-OSDet. This framework effectively fuses features from both modalities, alleviates cross-modal domain gaps, and results in significant improvements in detection performance.
The main contributions of this work are summarized as follows:
  • A Novel Optical-SAR Dataset: M4-SAR is a novel optical-SAR fusion object detection dataset with multi-resolution, multi-polarization, and multi-scene characteristics, containing 112,174 image pairs and 981,862 instances. This dataset comprehensively covers complex scenes and diverse target categories frequently encountered in multi-source fusion detection tasks, thereby providing a robust foundation for advancing research in this field.
  • Multi-source Evaluation Toolkit (MSRODet) and E2E-OSDet: We construct MSRODet, a comprehensive benchmarking toolkit incorporating state-of-the-art fusion object detection methods tailored for optical-SAR fusion object detection, which provides a standardized and reproducible evaluation platform based on the M4-SAR dataset. In addition, we propose E2E-OSDet, a novel end-to-end optical-SAR fusion object detection framework, which effectively mitigates significant cross-domain discrepancies between optical and SAR modalities at the feature level.
@@
2 M4-SAR Dataset Construction
2.1 Overview of the Proposed Dataset Construction Pipeline
@@
Fig. 2 illustrates the overall pipeline of the proposed dataset construction process. In the first stage, we acquire optical and SAR images from Sentinel-1 and Sentinel-2 satellites provided by the European Space Agency (ESA). In the second stage, the acquired images are pre-processed and aligned. After completing the alignment, we crop aligned image pairs into fixed-size patches to generate initial unlabeled samples. In the third stage, we introduce a semi-supervised annotation strategy guided by optical images to ensure high-quality labeling.
2.2 Data Collection
@@
As shown in Fig. 2 (a), we collected the required data via the open-access platform provided by the ESA. A map-based selection tool was used to specify areas of interest, covering representative coastal cities, including Tianjin, Los Angeles, Hong Kong, and London. We then retrieved SAR images along with their corresponding digital elevation model (DEM) data according to specific criteria. For optical imagery, Sentinel-2 images covering the same regions as the Sentinel-1 data were acquired, specifically including the red (R), green (G), and blue (B) spectral bands. To minimize variations caused by temporal discrepancies between the two modalities, we selected image pairs with a time interval of less than 10 days. This setting ensures that targets of interest (e.g., bridges, oil tanks) maintain structural consistency within the short interval, thereby enhancing the accuracy of subsequent instance-level alignment. The product type for optical images is L2A, with cloud coverage limited to 0–10% (the corresponding cloud-free (0%) images are used for high-quality annotation) to guarantee high visual fidelity. The SAR images are provided in the Single Look Complex (SLC) format to preserve full coherence information. The collected data span the period from 2020 to 2022 and encompass multiple spatial resolutions and polarization modes. The optical images have 10m, 20m, and 60m resolutions, and the SAR data cover both vertical-horizontal (VH) and vertical-vertical (VV) polarizations.
2.3 Data Processing